OpenStudy (anonymous):

solve using Cramer's Rule only : 3x+4y-3z=5 3x-2y+4z=7 3x+2y-z=3

OpenStudy (anonymous):

The rule is described here http://www.purplemath.com/modules/cramers.htm

OpenStudy (anonymous):

The system can be written in the form of a matrix like this: \[\left[\begin{matrix}3 & 4 & 5 \\ 3 & -2 & 4 \\ 3 & 2 & -1\end{matrix}\right] \left[\begin{matrix}5 \\ 7 \\3\end{matrix}\right]\] Confirm if this is understandable

OpenStudy (anonymous):

yes can you please show me how to solve it step by step plz

OpenStudy (anonymous):

hello, yes So cramer's rule is being used to efficiently solve for x, y or z which one do you want to solve ?

OpenStudy (anonymous):

x y and z

OpenStudy (anonymous):

please and thank you

OpenStudy (anonymous):

ok lets begin by solving X so to use cramer's rule here, we simply replace all entries belonging to x variable with the solution vector. The first entry in each row is the one that is connected to x